OpenStudy (anonymous):

what is the time taken by the piston to reach tdc from bdc...... guys please i need it urgent

OpenStudy (anonymous):

This is the cycle time. Let's take a 4-cycle engine as an example. Each of the four cycles (intake, compression, expansion (power), and exhaust) is defined as the piston moving from BDC to TDC or TDC to BDC.

OpenStudy (anonymous):

k mate.. im asking about the approximate time taken by the piston to complete one cycle say intake (what is the time taken by the piston to reach bdc from tdc in int intake process)...please mate i needit urgent

OpenStudy (anonymous):

The cycle time will depend on the engine speed. Let's assume that engine speed is expressed in RPM's (revolutions per minute) and that each revolution is composed of two cycles. Through dimensional analysis, we can find a conversion from RMP's to \(\rm sec/ cycle\) to be \[{\rm \sec \over cycle}~~ \alpha {\rm 1 \over RMP} ={\rm \min \over rev} \cdot {60 \sec \over \min} \cdot {\rm rev \over 2 cycle} = {\rm 30 \sec \over cycle}\] Therefore, if the engine is operating at 3,000 RPM, the sec/cycle will be equal to \[{\rm \sec \over cycle} = {30 \over 3000} = 0.01 \sec = 10 \rm ms\] If the engine is operating at 1,500 RPM, \[{\rm \sec \over cycle} = {\rm 30 \over 1500} = 0.02 \sec = \rm 20 ms\]

OpenStudy (anonymous):

thank u mate .....can u please tell me the impact force on the piston that the combusted fuel will cause

OpenStudy (anonymous):

That force is hard to determine. Considering that combustion events are still not well know and all equations relating to them are phenomenological, it is hard to determine this force based on engine parameters. We can determine this force from a torque measurement on a dynamometer. However, this is still tricky. If we know the torque, the piston diameter (bore), the radius of the crank shaft, and the angle of the crank shaft expressed in "aTDC" (after top dead center), we might be able to estimate the force exerted on the piston face. If we have a pressure vs. engine position graph, this is easy. \[F_p = P A_p\] We can use the maximum pressure to determine the maximum force. We can develop a kinetic equation if we measure the torque of the engine using a dynamometer but lack a pressure vs. engine position graph. \[I_e \ddot \theta = F_p A_p P r_{CS} - \tau\]Note that the dynamometer will measure maximum torque when \(I \ddot \theta = 0\), therefore we get\[F_P = {\tau \over A_p P r_{CS}}\]